# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
For the second term on the right-hand side of (7.15), note that $\mathsf{E}[(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}]\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}=0$ for all $i,j\in\{1,\ldots,n\}$, so that (7.10) implies $\mathsf{E}[\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})\mid\boldsymbol{X}_{1:n}]=0$, and thus
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi_{i}(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]$
$\quad=\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi_{i}(\boldsymbol{X}_{i})}\mathsf{E}\left[\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}[\mathds{1}_{\{Y_{i}\leq k/m\}}\mid\boldsymbol{X}_{1:n}]\,b_{m,k}(y)\right]$
@@
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]=\mathcal{O}(n^{-1}).$
@@
Let $y\in(0,1)$. We start by looking at the second term in the decomposition (7.9) and we decompose it as in (7.15). The first term on the right-hand side of (7.15) is $\mathcal{O}_{L^{2}}(n^{-1/2})$:
$\mathsf{E}\left[\left(n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right)^{2}\right]$
$\quad\leq n^{-2}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\mathsf{E}\left[\left|\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}\right|\right$.
$\hskip 113.81102pt\left.\left|\frac{(\delta_{j}-\pi(\boldsymbol{X}_{j}))}{\pi_{j}(\boldsymbol{X}_{j})^{2}}(\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi_{j}(\boldsymbol{X}_{j}))\mathds{1}_{\{Y_{j}\leq\ell/m\}}\right|\right]b_{m,k}(y)b_{m,\ell}(y)$
$\quad\ll\frac{n^{-2}}{\pi_{\min}^{4}}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\mathsf{E}\left[\left|\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})\right|\left|\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi_{j}(\boldsymbol{X}_{j})\right|\right]b_{m,k}(y)b_{m,\ell}(y)$
$\quad\ll\frac{n^{-2}}{\pi_{\min}^{4}}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\sqrt{\mathsf{E}\left[\left|\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})\right|^{2}\right]}\sqrt{\mathsf{E}\left[\left|\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi_{j}(\boldsymbol{X}_{j})\right|^{2}\right]}b_{m,k}(y)b_{m,\ell}(y)$
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi_{i}(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\quad=n^{-2}\sum_{k=0}^{m}\sum_{i,j=1}^{n}\frac{(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\frac{\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\qquad+n^{-2}\sum_{\ell=1}^{\infty}\sum_{k=0}^{m}\sum_{i,j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}\frac{(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
Combining (7.11) and (7.12) in the proof of Proposition 6, the second term on the right-hand side is $\mathcal{O}_{L^{2}}(n^{-1})$. The first term on the right-hand side is equal to
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1}),$
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.16)
$\quad=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1}).$
@@
$\widehat{F}_{n,m}(y)$ $=\widetilde{F}_{n,m}(y)-n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1})$
$=n^{-1}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-n^{-1}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)$
@@
$\mathsf{Var}(B_{n,m})$ $=n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\mathsf{Var}\left(\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\mid\boldsymbol{X}_{i}\right)\right]$
$\quad+n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\mathsf{E}\left[\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\mid\boldsymbol{X}_{i}\right]\right)$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{i})}{\pi_{i}(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)^{2}\right]$
$\quad+n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\frac{(\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right).$
Given that $\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{i})=\pi_{i}(\boldsymbol{X}_{i})(1-\pi_{i}(\boldsymbol{X}_{i}))$ and $\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi_{i}(\boldsymbol{X}_{i})=0$, it follows that
$\mathsf{Var}(B_{n,m})=n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)b_{m,k}(y)\right)^{2}\right].$
@@
$\mathsf{E}[B_{n,m}]$ $=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\right]b_{m,k}(y)$ (7.18)
$=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\mathsf{E}\left[\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\mid\boldsymbol{X}_{i}\right]\right]b_{m,k}(y)$
$=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\right]b_{m,k}(y)$
@@
$\mathsf{E}[A_{n,m}B_{n,m}]=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\delta_{i}(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right)\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)\right].$
@@
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{E}[\delta_{i}(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{i}]}{\pi_{i}(\boldsymbol{X}_{i})^{2}}\mathsf{E}\left[\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\mid\boldsymbol{X}_{i}\right]\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)\right]$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{E}[\delta_{i}(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{i}]}{\pi_{i}(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)^{2}\right].$
Since $\delta_{i}^{2}=\delta_{i}$ and $\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]=\pi_{i}(\boldsymbol{X}_{i})$, we have
$\mathsf{E}[A_{n,m}B_{n,m}]$ $=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\pi_{i}(\boldsymbol{X}_{i})(1-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y)\right)^{2}\right]$
$=n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)b_{m,k}(y)\right)^{2}\right]$
@@
$=\mathsf{Var}(\widetilde{F}_{n,m}(y))-n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)b_{m,k}(y)\right)^{2}\right].$

### v2

@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
For the second term on the right-hand side of (7.15), note that $\mathsf{E}[(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}]\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}=0$ for all $i,j\in\{1,\ldots,n\}$, so that (7.10) implies $\mathsf{E}[\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})\mid\boldsymbol{X}_{1:n}]=0$, and thus
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]$
$\quad=\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi(\boldsymbol{X}_{i})}\mathsf{E}\left[\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}[\mathds{1}_{\{Y_{i}\leq k/m\}}\mid\boldsymbol{X}_{1:n}]\,b_{m,k}(y)\right]$
@@
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]=\mathcal{O}(n^{-1}).$
@@
Let $y\in(0,1)$ be given. We start by looking at the second term on the right-hand side of (7.9) and we decompose it as in (7.15). The first term on the right-hand side of (7.15) is $\mathcal{O}_{L^{2}}(n^{-1/2})$:
$\mathsf{E}\left[\left(n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right)^{2}\right]$
$\quad\leq n^{-2}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\mathsf{E}\left[\left|\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}\right|\right$.
$\hskip 113.81102pt\left.\left|\frac{(\delta_{j}-\pi(\boldsymbol{X}_{j}))}{\pi(\boldsymbol{X}_{j})^{2}}(\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{j}))\mathds{1}_{\{Y_{j}\leq\ell/m\}}\right|\right]b_{m,k}(y)\,b_{m,\ell}(y)$
$\quad\ll\frac{n^{-2}}{\pi_{\min}^{4}}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\mathsf{E}\left[\left|\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})\right|\left|\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{j})\right|\right]b_{m,k}(y)\,b_{m,\ell}(y)$
$\quad\ll\frac{n^{-2}}{\pi_{\min}^{4}}\sum_{k,\ell=0}^{m}\sum_{i,j=1}^{n}\sqrt{\mathsf{E}\left[\left|\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})\right|^{2}\right]}\sqrt{\mathsf{E}\left[\left|\hat{\pi}_{j}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{j})\right|^{2}\right]}b_{m,k}(y)\,b_{m,\ell}(y)$
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\quad=n^{-2}\sum_{k=0}^{m}\sum_{i,j=1}^{n}\frac{(\delta_{j}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\frac{\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\qquad+n^{-2}\sum_{\ell=1}^{\infty}\sum_{k=0}^{m}\sum_{i,j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}\frac{(\delta_{j}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
By combining (7.11) and (7.12) in the proof of Proposition 6, the second term on the right-hand side is $\mathcal{O}_{L^{2}}(n^{-1})$. The first term on the right-hand side is equal to
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1}),$
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.16)
$\quad=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1}).$
@@
$\widehat{F}_{n,m}(y)$ $=\widetilde{F}_{n,m}(y)-n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)+\mathcal{O}_{L^{2}}(n^{-1})$
$=n^{-1}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-n^{-1}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)$
@@
$\mathsf{Var}(B_{n,m})$ $=n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\mathsf{Var}\left(\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\mid\boldsymbol{X}_{i}\right)\right]$
$\quad+n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\mid\boldsymbol{X}_{i}\right]\right)$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{i})}{\pi(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)^{2}\right]$
$\quad+n^{-2}\sum_{i=1}^{n}\mathsf{Var}\left(\frac{(\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right).$
Given that $\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{i})=\pi(\boldsymbol{X}_{i})(1-\pi(\boldsymbol{X}_{i}))$ and $\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi(\boldsymbol{X}_{i})=0$, it follows that
$\mathsf{Var}(B_{n,m})=n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)\,b_{m,k}(y)\right)^{2}\right].$
@@
$\mathsf{E}[B_{n,m}]$ $=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\right]b_{m,k}(y)$ (7.18)
$=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\mid\boldsymbol{X}_{i}\right]\right]b_{m,k}(y)$
$=n^{-1}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\right]b_{m,k}(y)$
@@
$\mathsf{E}[A_{n,m}B_{n,m}]=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\delta_{i}(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right)\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)\right].$
@@
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{E}[\delta_{i}(\delta_{i}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{i}]}{\pi(\boldsymbol{X}_{i})^{2}}\mathsf{E}\left[\sum_{k=0}^{m}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\mid\boldsymbol{X}_{i}\right]\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)\right]$
$=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\mathsf{E}[\delta_{i}(\delta_{i}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{i}]}{\pi(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)^{2}\right].$
Since $\delta_{i}^{2}=\delta_{i}$ and $\mathsf{E}[\delta_{i}\mid\boldsymbol{X}_{i}]=\pi(\boldsymbol{X}_{i})$, we have
$\mathsf{E}[A_{n,m}B_{n,m}]$ $=n^{-2}\sum_{i=1}^{n}\mathsf{E}\left[\frac{\pi(\boldsymbol{X}_{i})(1-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}\left(\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y)\right)^{2}\right]$
$=n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)\,b_{m,k}(y)\right)^{2}\right]$
@@
$=\mathsf{Var}(\widetilde{F}_{n,m}(y))-n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)\,b_{m,k}(y)\right)^{2}\right].$